Dimethylamine, (CH3)2NH, is commonly found in plants and food (in very small amounts). A solution of 0.100 M dimethylamine is 2.4% ionized. Calculate Kb for this base.
Answer #1

(CH3)2NH + H2O (CH3)2NH2+ + OH-

Let concentration of (CH3)2NH = c

degree of dissociation = x

Lets write (CH3)2NH as B and (CH3)2NH2+ as BH+

ICE table is

B BH+ OH-
Initial c 0 0
Change -cx +cx +cx
Equilibrium c - cx cx cx

Now, base dissociation constant

Kb =

Or, Kb =

Or, Kb =

Given, C = 0.100 M.

Percent ionization = 2.4%

So, x = (2.4/100) = 0.024

Then, Kb = [0.100× (0.024)2] /(1 - 0.024)]

= (5.76×10-5/0.976)

= 5.90× 10-5 .